# Personalized Layer Selection for Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) combine node attributes over a fixed granularity of the local graph structure around a node to predict its label. However, different nodes may relate to a node-level property with a different granularity of its local neighborhood, and using the same level of smoothing for all nodes can be detrimental to their classification. In this work, we challenge the common fact that a single GNN layer can classify all nodes of a graph by training GNNs with a distinct personalized layer for each node. Inspired by metric learning, we propose a novel algorithm, MetSelect [^1] Due to IP restrictions, we are still working on getting the code approval and plan to release it upon acceptance, to select the optimal representation layer to classify each node. In particular, we identify a prototype representation of each class in a transformed GNN layer and then, classify using the layer where the distance is smallest to a class prototype after normalizing with that layer’s variance. Results on $10$ datasets and $3$ different GNNs show that we significantly improve the node classification accuracy of GNNs in a plug-and-play manner. We also find that using variable layers for prediction enables GNNs to be deeper and more robust to poisoning attacks. We hope this work can inspire future works to learn more adaptive and personalized graph representations.
@@ 1 Introduction
Figure 1: GNNs with personalized prediction layer can empower existing GNNs for more accurate, personalized, robust, and deeper node classification (e.g., classifying users as conservative (purple) or liberal (green)). For example, while some nodes benefit from their neighbors, others may be better off by predicting using only their attributes.
@@ 2 Related Work
Layer aggregation methods such as JK-net, try to learn personalized representations by aggregating different representations through pooling operations and/or skip connections for node classification (Xu et al., 2018b; Dwivedi et al., 2023; Rusch et al., 2022). However, they do not adaptively select a single layer for classifying a given node but rather merge the representations for all nodes in the same way. This reduces the interepretability and increases the complexity of the underlying network by predicting through non-linear pooled transformation of the GNN representations. On the other hand, simpler layer selection of MetSelect preserves the interpretability and the analytic simplicity of the underlying GNN as the predictions are still made through the GNN embeddings.
@@
Based on the observations that a GCN can only approximate a low-pass filter (Nt & Maehara, 2019; Di Giovanni et al., 2022; Yan et al., 2022), new models have been proposed to address the heterophilic (Pei et al., 2020; Luan et al., 2022; Eliasof et al., 2023; Chien et al., 2020) and closely-related oversmoothing (Bodnar et al., 2022; Yan et al., 2022) challenge. However, these models achieve these enhancements by modifying the underlying message-passing architecture, while we present a plug-and-play approach to show that these issues can possibly be addressed by adaptive layer selection. NDLS (Zhang et al., 2021) also finds node-specific smoothing level, their method is based on a simple feature smoothing model and is agnostic of the underlying GNN and the label distribution. Furthermore, while they hypothesize that the optimal layer is the one at which the influence of other nodes is similar to the stationary influence, we hypothesize that given a GNN, it is the one that minimizes the distance from the class prototype in the corresponding embedding space.
@@
Certain methods aim to identify useful neighbors for each node separately to better classify them. Self-interpretable (Dai & Wang, 2021; Zhang et al., 2022; Feng et al., 2022) and attention-based (Veličković et al., 2017; Shi et al., 2020; Brody et al., 2021) graph neural networks weigh nodes based on the similarity between different nodes in the whole graph or the local graph structure. While this implicitly weighs each neighbor differently, the representations of all nodes are still formed from a fixed number of message-passing steps. Thus, these are still susceptible to oversmoothing and are not scalable as opposed to their graph neural network counterparts. Instead, we provide plug-and-play improvements to out-of-the-shelf GNNs by simply presenting a way to select optimal smoothing level for each node.
@@ 4 Problem
(GNN with node-optimal prediction layer) Given a graph $\mathbf{G}=(\mathbf{X},(\mathcal{V},\mathcal{E}))$ and a GNN encoder $\mathbf{h}_{\Theta}(\cdot)$ of depth $L$, the objective is to identify a personalized layer $l^{*}(v)\in[0,L]$ for each node $v\in\mathcal{V}$ such that $\mathbf{h}^{(l^{*}(v))}(v)$ best predicts the true label $y(v)$.
As detailed in Section 1, this problem is important to study since it allows us to enhance the GNNs.
@@ 5 Methodology
As noted in Equation 1, each layer of GNN $\mathbf{h}^{(l)}$ represents the nodes in distinct ways (typically, using the $l$-hop neighborhood) Thus, one can separate the nodes based on their labels in each of these spaces. Then, we propose to find the best layer $l^{*}(v)$ for a node based on how confidently that layer predicts its class. To compare the confidence of the prediction from different representation spaces, we adopt distance-based class decoding and use the distance as the notion of the confidence (Salakhutdinov & Hinton, 2007). However, distance-based decoding can distort the softmax decoding that we do to predict the labels from these layers. Thus, we propose to use a transformed representation $\tilde{\mathbf{h}}^{(l)}=\mathbf{W}\mathbf{h}^{(l)}+\mathbf{b}$ to find the best layer $l^{*}$.
@@
Distance-based classifiers Rippel et al. (2015); Deng et al. (2019; 2020); Salakhutdinov & Hinton (2007) typically finds a prototype embedding of each class and then classifies the input to the class whose prototype is the closest in the embedding space. For instance, for layer $l$, the prototype of each class can simply be given as the mean embedding of the training nodes belonging to that class, i.e.,
@@
Time Complexity. The total training time taken per epoch by Algorithm 1 includes the time for (1) Finding the moments of each class cluster, (2) Finding the optimal layer for each node, (3) Calculating the loss over the training nodes, and (4) Updating the parameters using gradient descent. One can find the means $\bm{\mu}_{c}^{(l)}$ in time $O(L|\mathcal{Y}||\mathcal{V}_{tr}|d)$ and the variances $\mathbf{K}_{l}$ in time $O(L|\mathcal{V}_{tr}|d^{2})$, where $d$ denotes the embedding dimension. The optimal layers can be found using $O(|\mathcal{V}_{tr}|L|\mathcal{Y}|d^{2})$. One can then decode the probabilities and accumulate the loss in time $O(|\mathcal{V}_{tr}||\mathcal{Y}|d)$. This makes up $O(|\mathcal{V}_{tr}|L|\mathcal{Y}|d^{2}))$ while the forward pass through the GNN takes at least $O(Ld^{2})$ (due to multiplication with $d\times d$ weight matrices). Thus, MetSelect remains efficient as long as $|\mathcal{V}_{tr}|$ and $|\mathcal{Y}|$ are small. While we empirically find the running time to be similar, specific sampling techniques to reduce the graph size can also be used to scale to larger graphs (Chiang et al., 2019).
Space Complexity. As shown in Algorithm 1, we calculate the means and variance by updating them in an online manner over each node.
@@
9:        $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}(v_{i},y_{i};l^{*})$ [Eqns. 4, 6]
@@ 7 Results
7.1 Can MetSelect be used to enhance the performance in a plug-and-play manner?
Here, we test the impact of using personalized prediction layer to classify the nodes from different GNN representations. Table 2 reports the performance of different layer selection methods on different GNNs and datasets. One can note that our proposed MetSelect method almost always gives the best accuracy among the baselines. This can be observed from the lowest average difference from the maximum accuracy ($\leq$ 0.015). Standard GNNs that only use the final layer to classify all nodes are known to obtain good performance in homophilic datasets (Hu et al., 2020). In these cases, the three personalized layer selection methods (NDLSelect, AttnSelect, and MetSelect) preserve the performance of the test set across different GNNs for these datasets, with MetSelect being the most consistent among the three.
GNNs trained using MetSelect can ignore a node’s neighbors at a scope that is not useful for its classification. This explains the overall boost in performance on heterophilic datasets. Table 2 finds that MetSelect can indeed improve/match the performance of the GNNs with a positive increase in the average performance for GCN, GAT, and GIN. In particular, we find a constant improvement in the extremely label-heterophilic datasets such as Cornell, Texas, and Wisconsin. This is particularly because MetSelect learns to ignore the graph structure when predicting labels for many nodes since the graph structure is not aligned with the label distribution for these datasets. Note that this is identified by MetSelect in an automated manner without specifically encoding this preference in the method anywhere. Finally, we found that MetSelect provides the best overall performance being at least $4$ times closer to the best performance across datasets.
@@ 7.2 Does MetSelect enable deeper GCN models?
To test MetSelect’s effect at different depths, we use the distance-based loss function in Equation LABEL:eq:metselect_distance to efficiently train the model. Since we optimally choose the prediction layer to classify each node, we expect the performance of GCN+MetSelect models to preserve the test performance even when we increase the number of GNN layers. Figure 3 demonstrates this effect on Cora, Citeseer, and Pubmed datasets. We find that the accuracy of GCN+MetSelect drops by only up to $10$ accuracy points even for a depth as high as $128$, while the standard GCN model drops by at least $50$ points with just $32$ layers. This shows that MetSelect can indeed enable much deeper GCNs, thus, increasing its representative power. For ogbn-arxiv, we found that while GCN’s accuracy reduces from $0.67$ to $0.22$ and $0.06$ at a depth of $16$ and $32$, while MetSelect preserves the accuracy with a slight reduction from $0.66$ to $0.65$ and $0.62$.
7.3 Are the models trained using MetSelect more robust to poisoning attacks?
Untargeted poisoning attacks (Zügner et al., 2018; Sun et al., 2020) have demonstrated that the mean GNN performance can suffer heavily from a change in the graph structure during training. In Section 4, we have highlighted that MetSelect can potentially make the models more robust to such untargeted attacks by ignoring the neighborhood scope not useful for a node’s prediction and thus, potentially the perturbations. To test this, we use Mettack (Zügner et al., 2018) to perturb our training graph that is used to train a GCN model using the standard and the proposed MetSelect loss. In particular, we use the black-box version of Mettack, where it finds the attacks using a surrogate GCN model trained on the same dataset. These attacks are structural and represent the flip of a link, i.e., add the link if it doesn’t exist and delete otherwise. The total number of links flipped or the budget is fixed as a parameter, that we set to be $p\times|\mathcal{E}|$ where we vary $p\in[0,0.5]$ denoted by Mettack Link (%).
@@ 7.4 Analysis on MetSelect
We theoretically found the running time complexity of MetSelect to be higher than existing GNNs. However, we find empirically that it remains within an acceptable factor of $<2.5$ across different datasets to the original method. Table 7.4 shows the mean per-epoch time on the largest dataset (ogbn-arxiv) and the remaining small datasets. As compared to NDLSelect, we remain within reasonable bounds of $<1.5$, while sometimes doing even better. Thus, our method is equivalent to the one-time cost of training a proxy model as is done in NDLSelect.

### v2

@@ Abstract
Graph Neural Networks (GNNs) combine node attributes over a fixed granularity of the local graph structure around a node to predict its label. However, different nodes may relate to a node-level property with a different granularity of its local neighborhood, and using the same level of smoothing for all nodes can be detrimental to their classification. In this work, we challenge the common fact that a single GNN layer can classify all nodes of a graph by training GNNs with a distinct personalized layer for each node. Inspired by metric learning, we propose a novel algorithm, MetSelect, to select the optimal representation layer to classify each node. In particular, we identify a prototype representation of each class in a transformed GNN layer and then classify each node using the layer where the distance is smallest to a class prototype after normalizing with that layer’s variance. Results on $10$ datasets and $3$ different GNNs show that we significantly improve the node classification accuracy of GNNs in a plug-and-play manner. We also find that using variable layers for prediction enables GNNs to be deeper and more robust to poisoning attacks. We hope this work can inspire future works to learn more adaptive and personalized graph representations.
@@ 1 Introduction
Figure 1: GNNs with personalized prediction layer can empower existing GNNs for more accurate, personalized, robust, and deeper node classification (e.g., classifying unlabeled users (gray) as conservative (purple) or liberal (green)). For example, while some nodes benefit from their neighbors, others may be better off by predicting using only their attributes.
@@ 2 Related Work
Layer aggregation methods such as JK-net, try to learn personalized representations by aggregating different representations through pooling operations and/or skip connections for node classification (Xu et al., 2018b; Dwivedi et al., 2023; Rusch et al., 2022). However, they do not adaptively select a single layer for classifying a given node but rather merge the representations for all nodes in the same way. This reduces the interpretability and increases the complexity of the underlying network by predicting through non-linear pooled transformation of the GNN representations. On the other hand, the simpler layer selection of MetSelect preserves the interpretability and the analytic simplicity of the underlying GNN as the predictions are still made through the GNN embeddings. Furthermore, layer aggregation methods require all representations to lie in the same space and to have the same dimension while selecting optimal layers through MetSelect completely relaxes such assumptions.
@@
Based on the observations that a GCN can only approximate a low-pass filter (Nt & Maehara, 2019; Di Giovanni et al., 2022; Yan et al., 2022), new models have been proposed to address the heterophilic (Pei et al., 2020; Luan et al., 2022; Eliasof et al., 2023; Chien et al., 2020; Rossi et al., 2024) and closely-related oversmoothing (Bodnar et al., 2022; Yan et al., 2022) challenge. However, these models achieve these enhancements by modifying the underlying message-passing architecture, while we present a plug-and-play approach to show that these issues can possibly be addressed by adaptive layer selection. NDLS (Zhang et al., 2021) also finds node-specific smoothing level, their method is based on a simple feature smoothing model and is agnostic of the underlying GNN and the label distribution. Furthermore, while they hypothesize that the optimal layer is the one at which the influence of other nodes is similar to the stationary influence, we hypothesize that given a GNN, it is the one that minimizes the distance from the class prototype in the corresponding embedding space.
@@
Certain methods aim to identify useful neighbors for each node separately to better classify them. Self-interpretable (Dai & Wang, 2021; Zhang et al., 2022; Feng et al., 2022) and attention-based (Veličković et al., 2017; Shi et al., 2020; Brody et al., 2021) graph neural networks weigh nodes based on the similarity between different nodes in the whole graph or the local graph structure. While this implicitly weighs each neighbor differently, the representations of all nodes are still formed from a fixed number of message-passing steps. Thus, these are still susceptible to oversmoothing and are not scalable as opposed to their graph neural network counterparts. Instead, we provide plug-and-play improvements to out-of-the-shelf GNNs by simply presenting a way to select the optimal smoothing level for each node.
@@ 4 Problem
(Node-personalized layer selection in GNN) Given a graph $\mathbf{G}=(\mathbf{X},(\mathcal{V},\mathcal{E}))$ and a GNN encoder $\mathbf{h}_{\Theta}(\cdot)$ of depth $L$, the objective is to identify a personalized layer $l^{*}(v)\in[0,L]$ for each node $v\in\mathcal{V}$ such that $\mathbf{h}^{(l^{*}(v))}(v)$ best predicts the true label $y(v)$.
As detailed in Section 1, this problem of selecting representations is important to study since it allows us to
@@ 5 Methodology
As noted in Equation 1, each layer of GNN $\mathbf{h}^{(l)}$ represents the nodes in distinct ways (typically, using the $l$-hop neighborhood). Thus, one can separate the nodes based on their labels in each of these spaces. Then, we propose to find the best layer $l^{*}(v)$ for a node based on how confidently that layer predicts its class. To compare the confidence of the prediction from different representation spaces, we adopt distance-based class decoding and use the distance as the notion of the confidence (Salakhutdinov & Hinton, 2007). However, distance-based decoding can distort the softmax decoding that we do to predict the labels from these layers. Thus, we propose to use a transformed representation $\tilde{\mathbf{h}}^{(l)}=\mathbf{W}\mathbf{h}^{(l)}+\mathbf{b}$ to find the best layer $l^{*}$.
@@
Distance-based classifiers Rippel et al. (2015); Deng et al. (2019; 2020); Salakhutdinov & Hinton (2007) typically find a prototype embedding of each class and then classify the input to the class whose prototype is the closest in the embedding space. For instance, for layer $l$, the prototype of each class can simply be given as the mean embedding of the training nodes belonging to that class, i.e.,
@@
Since learning one decoder per layer can be hard especially for higher number of layers, we also propose to directly leverage the class moments as a distance-based classification. Here, we minimize the distance of a training node from the prototype of the correct class in the closest layer $l$ and maximize the distance from the prototypes of other classes in all layers. Assuming $\alpha=1$, we propose
$\mathcal{L}_{\text{distance}}(v_{i},y_{i})=\max(0,\alpha+d_{i,y_{i}}^{(l^{*}(v_{i}|y_{i}))}+\log{\sum_{l=0}^{L}\sum_{j\neq y_{i}}\exp(-d_{i,j}^{(l)})}),$
However, we only use $\mathcal{L}_{\text{CE}}$ in the experiments unless otherwise mentioned.
@@
Time Complexity. The total training time taken per epoch by Algorithm 1 includes the time for (1) Finding the moments of each class cluster, (2) Finding the optimal layer for each node, (3) Calculating the loss over the training nodes, and (4) Updating the parameters using gradient descent. One can find the means $\bm{\mu}_{c}^{(l)}$ in time $O(L|\mathcal{Y}||\mathcal{V}_{tr}|d)$ and the variances $\mathbf{K}_{l}$ in time $O(L|\mathcal{V}_{tr}|d^{2})$, where $d$ denotes the embedding dimension. The optimal layers can be found using $O(|\mathcal{V}_{tr}|L|\mathcal{Y}|d^{2})$. One can then decode the probabilities and accumulate the loss in time $O(|\mathcal{V}_{tr}||\mathcal{Y}|d)$. This makes up $O(|\mathcal{V}_{tr}|L|\mathcal{Y}|d^{2}))$ while the forward pass through the GNN takes at least $O(Ld^{2})$ (due to multiplication with $d\times d$ weight matrices). Thus, MetSelect remains efficient as long as $|\mathcal{V}_{tr}|$ and $|\mathcal{Y}|$ are small.
@@
9:        $\mathcal{L}\leftarrow\mathcal{L}+\mathcal{L}(v_{i},y_{i};l^{*})$ [Eqns. 5, 6].
@@
While we empirically find the running time to be similar, specific sampling techniques to reduce the graph size can also be used to scale to larger graphs (Chiang et al., 2019).
Space Complexity. As shown in Algorithm 1, we calculate the means and variance by updating them in an online manner over each node. Thus, the space overhead of MetSelect does not depend on the number of nodes as it only involves $O(L|\mathcal{Y}|d)$ to store the means and $O(Ld^{2})$ to store the variance.
@@ 7 Results
7.1 Can MetSelect be used to enhance the performance of GNNs in a plug-and-play manner?
Here, we test the impact of using a personalized prediction layer to classify the nodes from different GNN representations. Table 2 reports the performance of different layer selection methods on different GNNs and datasets. One can note that our proposed MetSelect method almost always gives the best accuracy among the baselines. This can be observed from the lowest average difference from the maximum accuracy ($\leq$ 0.015). Standard GNNs that only use the final layer to classify all nodes are known to obtain good performance in homophilic datasets (Hu et al., 2020). In these cases, the three personalized layer selection methods (NDLSelect, AttnSelect, and MetSelect) preserve the performance of the test set across different GNNs for these datasets, with MetSelect being the most consistent among the three.
GNNs trained using MetSelect can ignore a node’s neighbors at a scope that is not useful for its classification. This explains the overall boost in performance on heterophilic datasets. Table 2 shows that MetSelect can indeed improve/match the performance of the GNNs with a positive increase in the average performance for GCN, GAT, and GIN. In particular, we find a constant improvement in the extremely label-heterophilic datasets such as Cornell, Texas, and Wisconsin. This is particularly because MetSelect learns to ignore the graph structure when predicting labels for many nodes since the graph structure is not aligned with the label distribution for these datasets. Note that this is identified by MetSelect in an automated manner without specifically encoding this preference in the method anywhere. We also note that performance on homophilic datasets does not show a significant increase, which can be explained due to the fact that MetSelect does not enhance the representation power of these models and selecting the most representative final layer tends to be the most powerful for these datasets. Overall, we find that MetSelect provides the best overall performance, being at least $4$ times closer to the best performance across datasets.
@@ 7.2 Does MetSelect enable deeper GCN models?
To test MetSelect’s effect at different depths, we use the distance-based loss function in Equation 6 to efficiently train the model at larger depths but find similar results to hold for the linear loss (i.e., Equation 6) till depth $L=8$. Since we optimally choose the prediction layer to classify each node, we expect the performance of GCN+MetSelect models to preserve the test performance even when we increase the number of GNN layers. Figure 3 demonstrates this effect on Cora, Citeseer, and Pubmed datasets. We find that the accuracy of GCN+MetSelect drops by only up to $10$ accuracy points even for a depth as high as $128$, while the standard GCN model drops by at least $50$ points with just $32$ layers. This shows that MetSelect can indeed enable much deeper GCNs, thus increasing its representative power. For ogbn-arxiv, we found that while GCN’s accuracy reduces from $0.67$ to $0.22$ and $0.06$ at a depth of $16$ and $32$, while MetSelect preserves the accuracy with a slight reduction from $0.66$ to $0.65$ and $0.62$.
7.3 Are the GNNs trained using MetSelect more robust to poisoning attacks?
Untargeted poisoning attacks (Zügner et al., 2018; Sun et al., 2020) have demonstrated that the mean GNN performance can suffer heavily from a change in the graph structure during training. In Section 4, we have highlighted that MetSelect can potentially make the models more robust to such untargeted attacks by ignoring the neighborhood scope not useful for a node’s prediction and, thus, potentially the perturbations. To test this, we use Mettack (Zügner et al., 2018) to perturb our training graph that is used to train a GCN model using the standard and the proposed MetSelect loss. In particular, we use the black-box version of Mettack, which finds the attacks using a surrogate GCN model trained on the same dataset. These attacks are structural and represent the flip of a link, i.e., add the link if it doesn’t exist and delete it otherwise. The total number of links flipped, or the budget is fixed as a parameter that we set to be $p\times|\mathcal{E}|$ where we vary $p\in[0,0.5]$ denoted by Mettack Link (%).
@@ 7.4 Analysis on MetSelect
Self-loops.
We further study the effect of including the feature information in the GNN and see if the gains remain consistent even in this setting. Particularly, we include the node features by adding a self-loop on each node which ensures that the node information is accounted for in all the aggregation layers (Hamilton et al., 2017). Table 4 shows that the performance remains consistently higher than the final layer selection baseline while highlighting the plug-and-play benefits of our method.
@@
We theoretically found the running time complexity of MetSelect to be higher than existing GNNs. However, we find empirically that it remains within an acceptable factor of $<2.5$ across different datasets to the original method. Table 3 shows the mean per-epoch time on the largest dataset (ogbn-arxiv) and the remaining small datasets. As compared to NDLSelect, we remain within reasonable bounds of $<1.5$, while sometimes doing even better. Thus, our method is equivalent to the one-time cost of training a proxy model as is done in NDLSelect.
@@
Table 4: Comparison of FinalSelect and MetSelect for GCN with self-loops.
FinalSelect $86.18_{1.15}$ $74.71_{1.66}$ $87.02_{0.45}$ $30.80_{0.92}$ $68.62_{1.81}$ $55.99_{1.02}$ $58.11_{9.56}$ $61.96_{7.23}$ $67.84_{5.76}$ $69.72$ $2.68$
MetSelect $83.72_{1.87}$ $73.05_{1.69}$ $86.03_{0.56}$ $30.99_{1.24}$ $63.75_{2.68}$ $51.22_{2.47}$ $66.22_{6.53}$ $76.67_{5.10}$ $71.62_{8.18}$ $66.77$ 1.77
Table 5: Comparison of MetSelect and layer aggregation methods in GCN using F1/100.
JKNet-Max 0.86 0.76 0.88 0.36 0.62 0.40 0.62 0.75 0.67 0.68 0.023
JKNet-Cat 0.86 0.76 0.88 0.35 0.61 0.42 0.66 0.78 0.71 0.69 0.011
JKNet-LSTM 0.71 0.72 0.88 0.35 0.50 0.29 0.70 0.79 0.75 0.68 0.046
JKNet-MetSelect 0.84 0.74 0.88 0.36 0.58 0.35 0.67 0.79 0.75 0.67 0.009
7.5 How does MetSelect compare with layer aggregation methods?
Since layer aggregation methods (Xu et al., 2018b) make the predictions from an aggregated/pooled layer using LSTM, max, or concatenation operators, they cannot solve Problem 2. In other words, they cannot select one of the existing GNN representations as a personalized layer to predict the node class. Thus, we avoid such a comparison for fairness since unlike layer aggregation methods, we do not enhance existing models’ representation power using non-linear activations.
For a fair comparison of the two methods, we create a setup where MetSelect also sees the set of pooled representations. Thus, we consider different layer aggregation strategies as multiple representations and select the best aggregation strategy among $\{\mathbf{h}^{(\text{LSTM})},\mathbf{h}^{(\text{CAT})},\mathbf{h}^{(\text{MAX})}\}$. We find in Table 5 that MetSelect can be used to select personalized aggregation layers for improved performance than using any of the aggregation. This shows complementary gains of our method in addition to simple GNN layer selection.